# Personalize Before Retrieve: LLM-based Personalized Query Expansion for User-Centric Retrieval

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) critically depends on effective query expansion to retrieve relevant information. However, existing expansion methods adopt uniform strategies that overlook user-specific semantics, ignoring individual expression styles, preferences, and historical context. In practice, identical queries in text can express vastly different intentions across users. This representational rigidity limits the ability of current RAG systems to generalize effectively in personalized settings. Specifically, we identify two core challenges for personalization: 1) user expression styles are inherently diverse, making it difficult for standard expansions to preserve personalized intent. 2) user corpora induce heterogeneous semantic structures—varying in topical focus and lexical organization—which hinders the effective anchoring of expanded queries within the user’s corpora space. To address these challenges, we propose Personalize Before Retrieve (PBR), a framework that incorporates user-specific signals into query expansion prior to retrieval. PBR consists of two components: P-PRF, which generates stylistically aligned pseudo feedback using user history for simulating user expression style, and P-Anchor, which performs graph-based structure alignment over user corpora to capture its structure. Together, they produce personalized query representations tailored for retrieval. Experiments on two personalized benchmarks show that PBR consistently outperforms strong baselines, with up to 10% gains on PersonaBench across retrievers. Our findings demonstrate the value of modeling personalization before retrieval to close the semantic gap in user-adaptive RAG systems. Our code is available at https://github.com/Zhang-Yingyi/PBR-code.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a pivotal paradigm for enhancing the capabilities of large language models (LLMs) (Lewis et al. 2020; Zhao et al. 2024; Zhang et al. 2025), leveraging a two-stage process: first retrieving relevant external corpus, then generating responses conditioned on the retrieved information (Cheng et al. 2023). The effectiveness of this paradigm hinges on the quality of query expansion by using LLM world knowledge (Culpepper et al. 2021; Song and Zheng 2024), which directly impacts the accuracy and relevance of the retrieval stage. Most current strategies generate pseudo answers (Gao et al. 2023) or candidate completions (Jia et al. 2023) using LLMs in a zero-shot or few-shot manner. These generated semantic representations, often combined with the original query, are used to retrieve information from a global external corpus.
Recently, with LLMs widely adopted as personalized assistants, many user queries are required to retrieve from individual histories and contexts, demanding retrieval mechanisms that go beyond generic corpora to leverage user-specific data. Existing query expansion strategies remain limited in their ability to capture user-specific semantics. In practice, the same textual query may convey vastly different intentions depending on a user’s preferences (Zou et al. 2023), background knowledge (Westhofen, Jung, and Neider 2025), or expression style (Neelakanteswara, Chaudhari, and Zamani 2024). As users exhibit diverse personas and contextual needs, it becomes increasingly important for query expansion methods to be personalized—capable of adapting to user-specific semantics within private corpora—so as to ensure accurate retrieval and enhance downstream generation.
@@
Motivated by these limitations, we aim to develop a personalized query expansion framework that adapts to individual user intent and expression as shown in Figure 1. However, introducing personalization into expansion brings two fundamental challenges: (1) User expression styles are inherently diverse. Users articulate intent using varied linguistic patterns—ranging from minimal prompts to elaborative reasoning—driven by personal habits, domain familiarity, or rhetorical preference (Neelakanteswara, Chaudhari, and Zamani 2024). These styles are often implicit and non-transferable, making it difficult to construct expansions that faithfully preserve user-specific semantics within a generalizable framework. (2) User corpora induce heterogeneous semantic structures. Users’ personal corpora often differ significantly in topical coverage, content organization, and linguistic granularity. This high degree of heterogeneity makes it difficult to locate reliable semantic regions related to the user query, and in turn, complicates the alignment of user-specific preferences. Without user-specific semantic anchors to guide expansion, queries are prone to drifting toward irrelevant regions, resulting in mismatched retrieval despite plausible surface semantics. These challenges raise a central question: How can we personalize query expansion—both in style and structure—before retrieval to align with individual user intent and corpus characteristics?
@@ 3.2 P-PRF
Generic query expansion methods often overlook the nuanced expression styles (Neelakanteswara, Chaudhari, and Zamani 2024) and implicit reasoning patterns behind user expression (Li et al. 2025), leading to mismatches in personalized retrieval. This challenge is exacerbated when queries are ambiguous or under-specified, requiring systems to infer intent beyond surface forms. Existing solutions typically apply uniform expansion strategies, failing to adapt to user-specific language signals.
@@ 3.3 P-Anchor
Existing methods are often designed to retrieve from a unified RAG corpus, lacking explicit grounding in the structural organization of individual user corpora (Tan et al. 2025; Wu et al. 2025). To address this, we propose P-Anchor, a structure-aware module that captures corpus-level preferences by identifying semantically central regions within a graph-structured user space.
@@
We represent the user’s prior corpus—such as interaction records with AI assistants—as a semantic graph $G=(\mathcal{C},\mathcal{E})$, following previous study (Tang et al. 2025), where $\mathcal{C}$ and $\mathcal{E}$ denote the sets of nodes and edges, respectively. The set of nodes $\mathcal{C}$ consists of history corpus vectors, where each node $\mathbf{c}_{i}$ as in Eq. (1) represents an encoded segment. The edges $\mathcal{E}$ are established based on pairwise similarity between the nodes. For each pair, we compute cosine similarity:
@@
We construct a sparse adjacency matrix $A$ to represent tht structural links within the user corpus, which retains only the top-$k_{2}$ neighbors exceeding a similarity threshold $\theta$:
@@
We following PageRank (Page et al. 1999; Haveliwala 2002) over $G$ to estimate a stationary distribution $\boldsymbol{\pi}$ reflecting node centrality. Let $\mathbf{P}=norm(\mathbf{A})$ be the row-normalized transition matrix, then:
@@ 3.4 PBR Fusion
To synthesize the signals captured in P-PRF and P-Anchor, we design a fusion mechanism that integrates stylistic pseudo-feedback with the query and user anchor. Recognizing that the utility of pseudo-feedback varies across users and queries—depending on individual expression patterns and the semantic openness of the query—we introduce a personalized and query-specific dynamic weight that balances the contribution of P-PRF according to its semantic proximity to both the original query $\mathbf{q}$ and the corpus anchor $\mathbf{c}_{\text{graph}}$. Specifically, we compute two weights $w_{1}$ and $w_{2}$ to quantify the semantic alignment of the pseudo-utterance and reasoning feedback, respectively:
@@
We retrieve from the user corpus $\mathcal{C}$ using faiss-based (Douze et al. 2024) nearest neighbor search with the fused query vector $\mathbf{q}^{*}$. This process leverages embedding-level alignment to recover user-relevant content.
@@ 4.1 Experiment Setting
We evaluate our method on two recent benchmarks tailored for personalized retrieval. PersonaBench (Tan et al. 2025) simulates user-specific queries over synthetic private data (e.g., preferences, behaviors, demographics), requiring models to align responses with personalized context. It includes both explicit and implicit user traits across diverse domains. LongMemEval (Wu et al. 2025) targets long-term interactive memory, where each user is associated with multi-turn dialogue traces. We use two subsets: LongMemEval-s (sparse history) and LongMemEval-m (dense history), to evaluate the effectiveness of retrieval under both short- and long-horizon personalization. More details are provided in Appendix 1.
@@
We compare our method with six strong baselines: Base, a static query retriever without expansion; HyDE (Gao et al. 2023), which generates hypothetical answers to guide retrieval; Query2Term (Jagerman et al. 2023), which performs keyphrase-based query expansion; MILL (Jia et al. 2023), which leverages multi-level logical structures; CoT (Wei et al. 2022), which applies chain-of-thought reasoning for better ranking; and ThinkQE (Lei, Shen, and Yates 2025).
@@
To verify the robustness of PBR, all methods are evaluated under three retrieval backbones: multi-qa-MiniLM-L6-cos-v1 (Wang et al. 2020), all-MiniLM-L6-v2 (Wang et al. 2020), and bge-base-en-v1.5 (Chen et al. 2024). For a fair comparison, all methods are using GPT-4o-mini (Achiam et al. 2023), and run 5 times. In PBR framework, we set $k_{1}=5$,$m=5$, $\theta=0.75$, and $k_{2}=10$ across Perosnabench and LongMemEval-s. The larger $k_{2}=50$ is particularly suited for LongMemEval-m, which contains richer user histories. More details are provided in Appendix 2.
@@ 5.1 Query Expansion in RAG System
Query expansion (QE) improves retrieval performance by simulating PRF. With the rise of LLMs, recent studies have explored LLMs as zero-shot QE engines. HyDE (Gao et al. 2023) and Query2Doc (Wang, Yang, and Wei 2023) demonstrate that prompting LLMs can yield precise expansions with PRF. Other works enhance semantic diversity or reasoning depth: MILL (Jia et al. 2023) employ mutual verification, and GRF (Mackie, Chatterjee, and Dalton 2023) apply LLM-based relevance feedback. ThinkQE (Lei, Shen, and Yates 2025) models expansion as a reasoning process, while LLMlingua (Jiang et al. 2023) focuses on compressing expansion prompts. GAR (Xia et al. 2024) further introduces a knowledge graph-aware QE framework.
@@ 5.2 Personalized RAG System
Personalized RAG system has recently emerged as a crucial direction for adapting LLMs to individual users. A recent survey by Li et al. (Li et al. 2025, 2023) provides a comprehensive overview, highlighting the importance of personalization in RAG. Approaches such as UniMS-RAG (Wang et al. 2024) integrate multi-source signals for dialogue personalization, while PersonaRAG (Zerhoudi and Granitzer 2024) enhances personalization via role-based agent frameworks. Salemi et al. (Salemi et al. 2024) further investigate LLM personalization by modeling language preferences and user goals. Complementary work like Wu et al. (Wu et al. 2024) explores the role of explicit user profiles, and Cohn et al. (Cohn et al. 2025) leverage interaction logs for educational agent adaptation. Domain-specific applications, such as personalized care systems (Yang et al. 2025) and memory-based assistant evaluation (Wu et al. 2025; Xu et al. 2025b, a), highlight the importance of long-term memory and evolving user states in effective RAG.
@@ 6 Conclusion
PersonaBench (Tan et al. 2025) [^1] https://github.com/SalesforceAIResearch/personabench is a benchmark dataset designed to evaluate personalized retrieval grounded in user-specific contexts. It includes six users, each associated with a heterogeneous personal corpus composed of (i) conversations with friends (Conv.), (ii) dialogues with AI assistants (AI.), and (iii) e-commerce purchase histories (E-com.). For each user, 40–50 short and ambiguous natural language queries are provided (e.g., “Do you know what my favorite color is?”), which require accurate grounding in the user’s prior behavior. This setup enables evaluation of models’ ability to align with diverse personal semantics. Table 4 summarizes the query and corpus statistics across users.
@@
LongMemEval (Wu et al. 2025) [^2] https://github.com/xiaowu0162/LongMemEval is a benchmark designed to evaluate the long-term memory capabilities of large language models, particularly their ability to retrieve task-relevant information from extended user-specific corpora. Each instance in the dataset consists of a factual question paired with a synthetic memory corpus simulating historical user data. The dataset includes two settings: LongMemEval-s, where each question is associated with approximately 50 memories, and LongMemEval-m, where each corpus contains over 500 memories. In total, both settings consist of 500 questions, allowing evaluation of memory retrieval precision under varying context lengths. This benchmark enables controlled analysis of personalized retrieval performance under realistic long-context scenarios.
@@
: All retrieval experiments are performed on a single NVIDIA A100 80GB GPU under Ubuntu 22.04. Query expansion is generated using GPT-4o-Mini with a temperature of 0. And each experiment we run 5 times.
@@
Our personalized query expansion framework (PBR) involves a two-stage prompt design to jointly simulate user-specific expression (P-PRF) and adaptively anchor reasoning paths (P-Anchor). Below we present the two corresponding prompts:
@@
These two prompts work in tandem: P-PRF synthesizes diverse user-style-aligned utterances, while P-Anchor explicitly grounds these expansions via structured reasoning. Together, they allow the model to generate personalized and semantically anchored queries before retrieval.
@@
To further evaluate the effect of personalized query expansion on generative outcomes, we conducted an additional experiment on the LongMemEval-S dataset under multi-qa-MiniLM-L6-cos-v1 retriever. Following prior work on RAG evaluation (Adlakha et al. 2024), we adopt a subspan Exact Match (EM) metric to account for partial correctness in generative responses. As shown in Table 8, PBR outperforms baseline query expansion approaches such as HyDE and CoT. Notably, PBR achieves the highest EM score of 42.6%, surpassing the non-personalized CoT (41.0%) and ThinkQE (41.8%). These results demonstrate that integrating personalized query representation not only enhances retrieval accuracy but also yields measurable gains in final generation quality, confirming the cross-stage benefit of the proposed pre-retrieval personalization.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) critically depends on effective query expansion to retrieve relevant information. However, existing expansion methods adopt uniform strategies that overlook user-specific semantics, ignoring individual expression styles, preferences, and historical context. In practice, identical queries in text can express vastly different intentions across users. This representational rigidity limits the ability of current RAG systems to generalize effectively in personalized settings. Specifically, we identify two core challenges for personalization: 1) user expression styles are inherently diverse, making it difficult for standard expansions to preserve personalized intent. 2) user corpora induce heterogeneous semantic structures—varying in topical focus and lexical organization—which hinders the effective anchoring of expanded queries within the user’s corpora space. To address these challenges, we propose Personalize Before Retrieve (PBR), a framework that incorporates user-specific signals into query expansion prior to retrieval. PBR consists of two components: P-PRF, which generates stylistically aligned pseudo feedback using user history for simulating user expression style, and P-Anchor, which performs graph-based structure alignment over user corpora to capture its structure. Together, they produce personalized query representations tailored for retrieval. Experiments on two personalized benchmarks show that PBR consistently outperforms strong baselines, with up to 10% gains on PersonaBench across retrievers. Our findings demonstrate the value of modeling personalization before retrieval to close the semantic gap in user-adaptive RAG systems.
Codehttps://github.com/Applied-Machine-Learning-Lab/PBR-code
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a pivotal paradigm for enhancing the capabilities of large language models (LLMs) (lewis2020retrieval; zhao2024retrieval; zhang2025lsrp), leveraging a two-stage process: first retrieving relevant external corpus, then generating responses conditioned on the retrieved information (cheng2023lift). The effectiveness of this paradigm hinges on the quality of query expansion by using LLM world knowledge (culpepper2021topic; song2024survey), which directly impacts the accuracy and relevance of the retrieval stage. Most current strategies generate pseudo answers (gao2023precise) or candidate completions (jia2023mill) using LLMs in a zero-shot or few-shot manner. These generated semantic representations, often combined with the original query, are used to retrieve information from a global external corpus.
Recently, with LLMs widely adopted as personalized assistants, many user queries are required to retrieve from individual histories and contexts, demanding retrieval mechanisms that go beyond generic corpora to leverage user-specific data. Existing query expansion strategies remain limited in their ability to capture user-specific semantics. In practice, the same textual query may convey vastly different intentions depending on a user’s preferences (zou2023users), background knowledge (westhofen2025temporal), or expression style (neelakanteswara2024rags). As users exhibit diverse personas and contextual needs, it becomes increasingly important for query expansion methods to be personalized—capable of adapting to user-specific semantics within private corpora—so as to ensure accurate retrieval and enhance downstream generation.
@@
Motivated by these limitations, we aim to develop a personalized query expansion framework that adapts to individual user intent and expression as shown in Figure 1. However, introducing personalization into expansion brings two fundamental challenges: (1) User expression styles are inherently diverse. Users articulate intent using varied linguistic patterns—ranging from minimal prompts to elaborative reasoning—driven by personal habits, domain familiarity, or rhetorical preference (neelakanteswara2024rags). These styles are often implicit and non-transferable, making it difficult to construct expansions that faithfully preserve user-specific semantics within a generalizable framework. (2) User corpora induce heterogeneous semantic structures. Users’ personal corpora often differ significantly in topical coverage, content organization, and linguistic granularity. This high degree of heterogeneity makes it difficult to locate reliable semantic regions related to the user query, and in turn, complicates the alignment of user-specific preferences. Without user-specific semantic anchors to guide expansion, queries are prone to drifting toward irrelevant regions, resulting in mismatched retrieval despite plausible surface semantics. These challenges raise a central question: How can we personalize query expansion—both in style and structure—before retrieval to align with individual user intent and corpus characteristics?
@@ 3.2 P-PRF
Generic query expansion methods often overlook the nuanced expression styles (neelakanteswara2024rags) and implicit reasoning patterns behind user expression (li2025survey), leading to mismatches in personalized retrieval. This challenge is exacerbated when queries are ambiguous or under-specified, requiring systems to infer intent beyond surface forms. Existing solutions typically apply uniform expansion strategies, failing to adapt to user-specific language signals.
@@ 3.3 P-Anchor
Existing methods are often designed to retrieve from a unified RAG corpus, lacking explicit grounding in the structural organization of individual user corpora (tan2025personabench; wu2025longmemeval). To address this, we propose P-Anchor, a structure-aware module that captures corpus-level preferences by identifying semantically central regions within a graph-structured user space.
@@
We represent the user’s prior corpus—such as interaction records with AI assistants—as a semantic graph $G=(\mathcal{C},\mathcal{E})$, following previous study (tang2025llm4tag), where $\mathcal{C}$ and $\mathcal{E}$ denote the sets of nodes and edges, respectively. The set of nodes $\mathcal{C}$ consists of history corpus vectors, where each node $\mathbf{c}_{i}$ as in Eq. (1) represents an encoded segment. The edges $\mathcal{E}$ are established based on pairwise similarity between the nodes. For each pair, we compute cosine similarity:
@@
We construct a sparse adjacency matrix $A$ to represent the structural links within the user corpus, which retains only the top-$k_{2}$ neighbors exceeding a similarity threshold $\theta$:
@@
We following PageRank (page1999pagerank; haveliwala2002topic) over $G$ to estimate a stationary distribution $\boldsymbol{\pi}$ reflecting node centrality. Let $\mathbf{P}=norm(\mathbf{A})$ be the row-normalized transition matrix, then:
@@ 3.4 PBR Fusion
To synthesize the signals captured in P-PRF and P-Anchor, we design a fusion mechanism that integrates stylistic pseudo-feedback with the query and user anchor. Recognizing that the utility of pseudo-feedback varies across users and queries—depending on individual expression patterns and the semantic openness of the query—we introduce a personalized and query-specific dynamic weight that balances the contribution of P-PRF according to its semantic proximity to both the original query $\mathbf{q}$ and the corpus anchor $\mathbf{c}_{\text{Anchor}}$. Specifically, we compute two weights $w_{1}$ and $w_{2}$ to quantify the semantic alignment of the pseudo-utterance and reasoning feedback, respectively:
@@
We retrieve from the user corpus $\mathcal{C}$ using faiss-based (douze2024faiss) nearest neighbor search with the fused query vector $\mathbf{q}^{*}$. This process leverages embedding-level alignment to recover user-relevant content.
@@ 4.1 Experiment Setting
We evaluate our method on two recent benchmarks tailored for personalized retrieval. PersonaBench (tan2025personabench) simulates user-specific queries over synthetic private data (e.g., preferences, behaviors, demographics), requiring models to align responses with personalized context. It includes both explicit and implicit user traits across diverse domains. LongMemEval (wu2025longmemeval) targets long-term interactive memory, where each user is associated with multi-turn dialogue traces. We use two subsets: LongMemEval-s (sparse history) and LongMemEval-m (dense history), to evaluate the effectiveness of retrieval under both short- and long-horizon personalization. More details are provided in Appendix 1.
@@
We compare our method with six strong baselines: Base, a static query retriever without expansion; HyDE (gao2023precise), which generates hypothetical answers to guide retrieval; Query2Term (jagerman2023query), which performs keyphrase-based query expansion; MILL (jia2023mill), which leverages multi-level logical structures; CoT (wei2022chain), which applies chain-of-thought reasoning for better ranking; and ThinkQE (lei2025thinkqe).
@@
To verify the robustness of PBR, all methods are evaluated under three retrieval backbones: multi-qa-MiniLM-L6-cos-v1 (wang2020minilm), all-MiniLM-L6-v2 (wang2020minilm), and bge-base-en-v1.5 (chen2024bge). For a fair comparison, all methods use an advanced LLM and run 5 times. In PBR framework, we set $k_{1}=5$,$m=5$, $\theta=0.75$, and $k_{2}=10$ across Perosnabench and LongMemEval-s. The larger $k_{2}=50$ is particularly suited for LongMemEval-m, which contains richer user histories. More details are provided in Appendix 2.
@@ 5.1 Query Expansion in RAG System
Query expansion (QE) improves retrieval performance by simulating PRF. With the rise of LLMs, recent studies have explored LLMs as zero-shot QE engines. HyDE (gao2023precise) and Query2Doc (wang2023query2doc) demonstrate that prompting LLMs can yield precise expansions with PRF. Other works enhance semantic diversity or reasoning depth: MILL (jia2023mill) employ mutual verification, and GRF (mackie2023generative) apply LLM-based relevance feedback. ThinkQE (lei2025thinkqe) models expansion as a reasoning process, while LLMlingua (jiang2023llmlingua) focuses on compressing expansion prompts. GAR (xia2024knowledge) further introduces a knowledge graph-aware QE framework.
@@ 5.2 Personalized RAG System
Personalized RAG system has recently emerged as a crucial direction for adapting LLMs to individual users. A recent survey by Li et al. (li2025survey; li2023agent4ranking) provides a comprehensive overview, highlighting the importance of personalization in RAG. Approaches such as UniMS-RAG (wang2024unims) integrate multi-source signals for dialogue personalization, while PersonaRAG (zerhoudi2024personarag) enhances personalization via role-based agent frameworks. Salemi et al. (salemi2024lamp) further investigate LLM personalization by modeling language preferences and user goals. Complementary work like Wu et al. (wu2024understanding) explores the role of explicit user profiles, and Cohn et al. (cohn2025personalizing) leverage interaction logs for educational agent adaptation. Domain-specific applications, such as personalized care systems (yang2025improving) and memory-based assistant evaluation (wu2025longmemeval; xu2025towards; xu2025align), highlight the importance of long-term memory and evolving user states in effective RAG.
@@ 6 Conclusion
Acknowledgments
This research was supported by the National Natural Science Foundation of China (NSFC) under Grant 72071029, 72231010, 62502404, and the Graduate Research Fund of the School of Economics and Management of Dalian University of Technology (No. DUTSEMDRFKO1). This research was partially supported by Hong Kong Research Grants Council (Research Impact Fund No.R1015-23, Collaborative Research Fund No.C1043-24GF, General Research Fund No.11218325), Institute of Digital Medicine of City University of Hong Kong (No.9229503), and Huawei (Huawei Innovation Research Program).
@@ Acknowledgments
PersonaBench (tan2025personabench) [^1] https://github.com/SalesforceAIResearch/personabench is a benchmark dataset designed to evaluate personalized retrieval grounded in user-specific contexts. It includes six users, each associated with a heterogeneous personal corpus composed of (i) conversations with friends (Conv.), (ii) dialogues with AI assistants (AI.), and (iii) e-commerce purchase histories (E-com.). For each user, 40–50 short and ambiguous natural language queries are provided (e.g., “Do you know what my favorite color is?”), which require accurate grounding in the user’s prior behavior. This setup enables evaluation of models’ ability to align with diverse personal semantics. Table 4 summarizes the query and corpus statistics across users.
@@
LongMemEval (wu2025longmemeval) [^2] https://github.com/xiaowu0162/LongMemEval is a benchmark designed to evaluate the long-term memory capabilities of large language models, particularly their ability to retrieve task-relevant information from extended user-specific corpora. Each instance in the dataset consists of a factual question paired with a synthetic memory corpus simulating historical user data. The dataset includes two settings: LongMemEval-s, where each question is associated with approximately 50 memories, and LongMemEval-m, where each corpus contains over 500 memories. In total, both settings consist of 500 questions, allowing evaluation of memory retrieval precision under varying context lengths. This benchmark enables controlled analysis of personalized retrieval performance under realistic long-context scenarios.
@@
: All retrieval experiments are performed on an advanced GPU under Ubuntu 22.04. Query expansion is generated using an advanced LLM with a temperature of 0. And each experiment we run 5 times.
@@
Our P-PRF involves a two-stage prompt design to jointly simulate user-specific expression with utterance and reason.
@@
To further evaluate the effect of personalized query expansion on generative outcomes, we conducted an additional experiment on the LongMemEval-S dataset under multi-qa-MiniLM-L6-cos-v1 retriever. Following prior work on RAG evaluation (adlakha2024evaluating), we adopt a subspan Exact Match (EM) metric to account for partial correctness in generative responses. As shown in Table 8, PBR outperforms baseline query expansion approaches such as HyDE and CoT. Notably, PBR achieves the highest EM score of 42.6%, surpassing the non-personalized CoT (41.0%) and ThinkQE (41.8%). These results demonstrate that integrating personalized query representation not only enhances retrieval accuracy but also yields measurable gains in final generation quality, confirming the cross-stage benefit of the proposed pre-retrieval personalization.